The decomposition of SO2Cl2 is first order and has a rate constant of + 2.90 x 10–4 s–1. What is the concentration of SO2Cl2 after 840 seconds given that the initial concentration of SO2Cl2 is 0.0325M ? (hint: your answer should have three significant figures, not include units, and not be written in scientific notation)
Answer #1

Answer-

constant k is given by

-ln (1-XA)= kt      (1)

3) given CA =? and CAO= 0.0325 ,at t=840 s  

and 1-XA=CA/CAO =(CA =conc. at any time t,CAO=Initial concentration)

From (1) –ln (1-XA)= 2.90*10-4*840 =0.2436

1-XA= exp(-0.2436)= 1.27583

CA/CAO =1.27583 and so CA =1.27583*0.0325=0.04146